### v1

# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

Abstract.
Superconvergent and divergence-free finite element methods for the Stokes equation are developed. The velocity and pressure are discretized using $H(\mathrm{div})$-conforming vector elements and discontinuous piecewise polynomials. The discrete formulation employs a weak deviatoric gradient operator built with tangential-normal continuous finite elements for traceless tensors, requiring no stabilization. Optimal and superconvergent error estimates are established. The method connects to nonconforming virtual element and pseudostress-velocity-pressure mixed formulations. Numerical experiments verify the theory.
Key words and phrases:
Stokes equation, divergence-free, superconvergence, mixed finite element method
2020 Mathematics Subject Classification:
65N12; 65N22; 65N30;
The first author was supported by NSF DMS-2309777 and DMS-2309785. The second author was supported by the National Natural Science Foundation of China Project 12171300.
1. Introduction
In this work, we develop superconvergent and divergence-free finite element methods for the Stokes equation on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) with external force $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{d})$:
(1) $\begin{cases}-\Delta\boldsymbol{u}-\nabla p=\boldsymbol{f}&\text{in }\Omega,\\ \operatorname{div}\boldsymbol{u}=0&\text{in }\Omega,\\ \boldsymbol{u}=0&\text{on }\partial\Omega,\end{cases}$
where $\boldsymbol{u}$ and $p$ denote the velocity and pressure.
Classical stable finite element pairs for (1) include the Taylor–Hood element [50], the MINI element [3], and the nonconforming $P_{1}$$P_{0}$ element [25]. More examples and a detailed discussion are given in [10]. However, these classical pairs are not pointwise divergence-free, which weakens mass conservation and reduces pressure robustness in error estimates. A review of the importance of divergence-free constraints and their numerical implications can be found in [41].
Divergence-free finite element pairs include the Scott–Vogelius element [48], which requires special mesh conditions; smooth finite element pairs [56, 17, 18] with supersmooth degrees of freedom; conforming pairs on split meshes [6, 55, 22, 28, 32, 38]; conforming rational pairs [33, 34]; and low-order, normal-continuous nonconforming pairs [43, 49, 54]. We also refer to [8, 52, 51] for discontinuous Galerkin methods, and to [9, 20, 53, 40] for virtual element methods.
One can impose the divergence-free condition by using an $H(\operatorname{div})$-conforming finite element space for the velocity. Specifically, we use the Raviart–Thomas (RT) element space [47, 45] or the Brezzi–Douglas–Marini (BDM) element space [12, 46]:
$\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}:=\{\boldsymbol{v}_{h}\in H_{0}(\operatorname{div},\Omega):\boldsymbol{v}_{h}|_{T}\in\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x}\;\text{ for }T\in\mathcal{T}_{h}\}$
for the velocity, with integers $\ell=k$ (${\rm RT}_{k}$) or $\ell=k-1$ (${\rm BDM}_{k}$), and the discontinuous piecewise polynomial space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for the pressure.
Since $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\not\subset H^{1}(\Omega;\mathbb{R}^{d})$, the discretization of $\Delta\boldsymbol{u}$ is nontrivial. One approach is to use discontinuous Galerkin (DG) methods for the Laplacian operator, as in [52, 23]. The DG method in [52] introduces a penalty term to ensure stability, while the hybridizable DG (HDG) method in [23] employs $k$th-order polynomial spaces for all variables. In fact, it suffices to take $p_{h}\in\mathbb{P}_{k-1}(\mathcal{T}_{h})$ when the velocity $\boldsymbol{u}_{h}\in{\rm BDM}_{k}$.
In the vorticity–velocity–pressure formulation [26, 27], the vorticity $\omega={\rm curl\,}\boldsymbol{u}$ is introduced, and
$\Delta\boldsymbol{u}={\rm curl\,}\omega-{\rm grad\,}\operatorname{div}\boldsymbol{u}={\rm curl\,}{\rm curl\,}\boldsymbol{u}-{\rm grad\,}\operatorname{div}\boldsymbol{u}$
is interpreted as the Hodge Laplacian. However, as observed in [26, 27], for no-slip Dirichlet boundary conditions ($\boldsymbol{u}=0$ on $\partial\Omega$), there are stability issues and reduced convergence rates. This was later explained in [4]: the suboptimal convergence rate comes from the lack of a proper Hilbert complex in this setting.
The lost convergence rate can be partly recovered by mesh symmetry. In the triangular MAC (TMAC) schemes for the Stokes equation [21], mass lumping is applied to the vorticity so that it can be eliminated, and the velocity and pressure convergence rates are improved by mesh symmetry. Using the element pair $\textrm{RT}_{0}$$\textrm{P}_{0}$ for velocity and pressure (i.e., $k=\ell=0$) yields
$|I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{\min\{1,s\}}|\ln h|^{1/2},$
where $s\geq\tfrac{1}{2}$ measures mesh symmetry and $I^{\rm\operatorname{div}}_{k,\ell}$ is the interpolation operator into $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$. With the pair $\textrm{BDM}_{1}$$\textrm{P}_{0}$ (i.e., $k=1$, $\ell=0$), the TMAC scheme achieves
$|I_{1,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{1+\min\{1/2,s\}}.$
In [30, 31], the mass-conserving mixed formulation with stresses (MCS) finite element method was developed for the pseudostress–velocity–pressure formulation, also employing $H(\operatorname{div})$-conforming velocity spaces. The stress space in [31] is slightly nonconforming, leading to the standard $h^{k}$ convergence instead of the superconvergent $h^{k+1}$ rate. In [30], a weak symmetry constraint was introduced and the stress space was enriched, yielding $h^{k+1}$ convergence for the $\mathrm{RT}_{k}$$\mathrm{P}_{k}$ pair with $k\geq 1$.
A divergence free weak virtual element method was developed in [20], which requires stabilization and does not yield superconvergence. The main idea is to introduce a piecewise vector-valued polynomial $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}:=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ defined on faces to enforce tangential continuity. We follow this approach and improve it to a stabilization-free scheme with superconvergence. Define the weak deviatoric gradient operator element-wise by
$(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu}),\boldsymbol{\tau})_{T}=-(\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v},\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}$
for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$. The tensor space is reduced to the traceless tensor $\mathbb{T}$, since for $\boldsymbol{\sigma}:={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{tr}{\rm grad\,}\boldsymbol{u}=\operatorname{div}\boldsymbol{u}=0$.
We propose the following mixed finite element method: find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}$, and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ with integers $k\geq 0$ and $\ell\in\{k,k-1\}$, such that
(2a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))+(\operatorname{div}\boldsymbol{v}_{h},p_{h})$ $=(\boldsymbol{f},\boldsymbol{v}_{h}),$
(2b) $(\operatorname{div}\boldsymbol{u}_{h},q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, and $q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$. Notably, the discrete method (2) requires no additional stabilization, despite relying on the weak differential operator $\operatorname{dev}{\rm grad\,}_{w}$.
We derive superconvergent estimates for the velocity and pressure:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|+\|Q_{\ell}p-p_{h}\|\lesssim h^{k+1}|\boldsymbol{\sigma}|_{k+1},$
where $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u},$ and $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$. Unlike the TMAC scheme [21], such superconvergence does not require the symmetry of the triangulation. It is further utilized to construct a new superconvergent approximation of $\boldsymbol{u}$.
Our method allows $\ell=0$ an important case not covered by MCS [30, 31]. We provide an illustration of the case $(k,\ell)=(0,0)$, which can be treated as a generalization of the classical MAC scheme [35] on rectangular grids to triangulations and achieve first order $h$ convergence, no loss of convergence compared with the vorticity-velocity-pressure formulation. One can use $\textrm{BDM}_{1}$$\textrm{P}_{0}$ pair, $(k,\ell)=(1,0)$, to get a second order $h^{2}$ scheme.
We propose a hybridization of (2) by relaxing the normal continuity of the velocity, which enables the choice $(k,\ell)=(0,-1)$. The velocity is approximated by the piecewise constant vector space, and the normal continuity is enforced through a nonconforming linear element space for the pressure $p_{h}$. Even in this low-order setting, the method achieves at least first-order convergence, which is a notable feature.
The remainder of this paper is organized as follows. Section 2 discusses the weak div stability and the distributional discretization of the vector Laplacian. Section 3 develops mixed finite element methods for the Stokes equation, and Section 4 presents their equivalent discrete formulations. Finally, Section 5 reports numerical experiments that confirm and illustrate the theoretical results.
2. Distributional Discretization of the Vector Laplacian
In this section, we discretize the weak divergence operator
$\operatorname{div}:L^{2}(\Omega;\mathbb{T})\to H^{-1}(\Omega;\mathbb{R}^{d}),$
and apply it to the vector Laplacian problem. We introduce weak divergence and deviatoric gradient operators, establish their inf-sup stability, and derive a superconvergent error estimate for divergence-free velocity spaces.
2.1. Notation
Let $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) be a bounded polytope with boundary $\partial\Omega$. For a bounded domain $D$ and integer $m\geq 0$, let $H^{m}(D)$ be the Sobolev space with norm $\|\cdot\|_{m,D}$ and semi-norm $|\cdot|_{m,D}$, and let $H_{0}^{m}(D)$ be the closure of $C_{0}^{\infty}(D)$ under $\|\cdot\|_{m,D}$. Set $L^{2}(D)=H^{0}(D)$ with inner product $(\cdot,\cdot)_{D}$, and write $\|\cdot\|_{0,D}$ as $\|\cdot\|_{D}$. When $D=\Omega$, the subscript $D$ is omitted. Let $h_{D}$ denote the diameter of $D$, and $\boldsymbol{n}_{\partial D}$ the unit outward normal, abbreviated as $\boldsymbol{n}$.
We use
$H(\operatorname{div},D)$ $=\{\boldsymbol{v}\in L^{2}(D;\mathbb{R}^{d}):\operatorname{div}\boldsymbol{v}\in L^{2}(D)\},$
$H_{0}(\operatorname{div},D)$ $=\{\boldsymbol{v}\in H(\operatorname{div},D):\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial D\}.$
Let $L_{0}^{2}(D)=\{q\in L^{2}(D):\int_{D}q\,\,{\rm d}x=0\}.$ For any $B\subseteq L^{2}(D)$, denote $B/\mathbb{R}=B\cap L_{0}^{2}(D)$.
Let $\{\mathcal{T}_{h}\}_{h>0}$ be a regular family of simplicial meshes of $\Omega$, where $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}=\operatorname{diam}(T)$. Denote by $\mathcal{F}_{h}$ and $\mathring{\mathcal{F}}_{h}$ the sets of all faces and interior faces, respectively. For each face $F\in\mathcal{F}_{h}$, fix a unit normal $\boldsymbol{n}_{F}$. For $T\in\mathcal{T}_{h}$, let $\mathcal{F}(T)$ be the set of all (d1)(d-1)-dimensional faces of $T$.
For a face $F$ and vector $\boldsymbol{v}\in\mathbb{R}^{d}$, define its tangential projection
$\Pi_{F}\boldsymbol{v}=(\boldsymbol{I}-\boldsymbol{n}_{F}\boldsymbol{n}_{F}^{\intercal})\boldsymbol{v}.$
For two adjacent elements $T^{\pm}$ sharing $F$, define the jump
$\llbracket v\rrbracket|_{F}=v^{+}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{+}}+v^{-}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{-}},$
and on a boundary face $F\subset\partial\Omega$, set $\llbracket v\rrbracket=v|_{F}\,\boldsymbol{n}_{F}\cdot\boldsymbol{n}$.
For a domain $D$ and integer $k\geq 0$, let $\mathbb{P}_{k}(D)$ be the space of polynomials on $D$ of degree $\leq k$, $\mathbb{H}_{k}(D)=\mathbb{P}_{k}(D)\setminus\mathbb{P}_{k-1}(D)$, and $Q_{k,D}$ the $L^{2}$-projection onto $\mathbb{P}_{k}(D)$. Set $\mathbb{M}=\mathbb{R}^{d\times d}$, $\mathbb{K}$ for skew-symmetric, and $\mathbb{T}$ for traceless matrices. For $\boldsymbol{\tau}\in\mathbb{M}$,
$\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\tfrac{1}{d}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I},\qquad\operatorname{skw}\boldsymbol{\tau}=\tfrac{1}{2}(\boldsymbol{\tau}-\boldsymbol{\tau}^{\intercal}).$
Define
$H^{s}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in H^{s}(T)\}=\prod_{T\in\mathcal{T}_{h}}H^{s}(T),$
$\mathbb{P}_{k}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in\mathbb{P}_{k}(T)\}=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T),$
$\mathbb{P}_{k}(\mathcal{F}_{h})=\{v\in L^{2}(\mathcal{F}_{h}):v|_{F}\in\mathbb{P}_{k}(F)\}=\prod_{F\in\mathcal{F}_{h}}\mathbb{P}_{k}(F),$
and similarly define $\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$ for interior faces only. For a space $B(D)$ or $B(\mathcal{T}_{h})$, define
$B(D;\mathbb{X})=B(D)\otimes\mathbb{X},\quad B(\mathcal{T}_{h};\mathbb{X})=B(\mathcal{T}_{h})\otimes\mathbb{X},$
where $\mathbb{X}\in\{\mathbb{R}^{d},\mathbb{M},\mathbb{T},\mathbb{K}\}$. Let $Q_{k}$ denote the $L^{2}$-projection operator from $L^{2}(\Omega)$ onto $\mathbb{P}_{k}(\mathcal{T}_{h})$; its vector- or tensor-valued version is also denoted by $Q_{k}$. We use ${\rm grad\,}_{h}$ ($\nabla_{h}$ for scalar function) and $\operatorname{div}_{h}$ to represent the element-wise gradient and divergence operators, respectively.
2.2. Weak div stability for the scalar Laplacian
We first consider the surjective map
$\operatorname{div}:L^{2}(\Omega;\mathbb{R}^{d})\to H^{-1}(\Omega),$
and discuss its distributional finite element discretization. This operator is part of the distributional de Rham complex introduced in [11].
The $L^{2}$ space is discretized by the discontinuous polynomial space
$\Sigma_{k}^{-1}(\mathcal{T}_{h};\mathbb{R}^{d})=\mathbb{P}_{k}(\mathcal{T}_{h};\mathbb{R}^{d}),$
which we abbreviate as $\Sigma_{k}^{-1}$. A natural norm on $\Sigma_{k}^{-1}$ is the standard $L^{2}$ norm $\|\cdot\|$.
Next, we introduce the product space of discontinuous polynomial spaces on $\mathcal{T}_{h}\times\mathcal{F}_{h}$:
$M_{k-1,k}^{-1}:=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathcal{F}_{h}).$
An element of $M_{k-1,k}^{-1}$ is denoted by $u=(u_{0},u_{b})$. A natural product-type inner product is defined by
$((u_{0},u_{b}),(v_{0},v_{b}))_{0,h}=(u_{0},v_{0})+\sum_{F\in\mathcal{F}_{h}}h_{F}(u_{b},v_{b})_{F},$
where the scaling factor $h_{F}$ ensures that the two terms are of comparable magnitude. The corresponding norm on $M_{k-1,k}^{-1}$ is
$\|u\|_{0,h}^{2}=\|u_{0}\|^{2}+\sum_{F\in\mathcal{F}_{h}}\|h_{F}^{1/2}u_{b}\|_{F}^{2}.$
The subspace
$\mathring{M}_{k-1,k}^{-1}=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$
can be viewed as a subspace of ${M}_{k-1,k}^{-1}$ by setting the values on boundary faces to zero.
For $\sigma\in\Sigma_{k}^{-1}$, define the operator $\operatorname{div}_{w}:\Sigma_{k}^{-1}\to\mathring{M}_{k-1,k}^{-1}$ by
$\operatorname{div}_{w}\sigma=\big(\operatorname{div}_{T}\sigma,-h_{F}^{-1}[\sigma\cdot\boldsymbol{n}]|_{F}\big)_{T\in\mathcal{T}_{h},~F\in\mathring{\mathcal{F}}_{h}}\in\mathring{M}_{k-1,k}^{-1}.$
When $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we have $\operatorname{div}_{w}\sigma=(\operatorname{div}\sigma,0)$ since $\sigma\cdot\boldsymbol{n}]|_{F}=0$ for all interior faces $F$; in this case we simply write $\operatorname{div}_{w}\sigma=\operatorname{div}\sigma$.
A graph norm associated with $\operatorname{div}_{w}$ is defined as
$\|\sigma\|_{\operatorname{div}_{w}}^{2}:=\|\sigma\|^{2}+\|\operatorname{div}_{w}\sigma\|_{0,h}^{2}=\|\sigma\|^{2}+\|\operatorname{div}_{h}\sigma\|^{2}+\|h^{-1/2}[\sigma\cdot\boldsymbol{n}]\|_{\mathring{\mathcal{F}}_{h}}^{2},$
where the scaling is chosen so that the last two terms are of comparable order. For $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we recover $\|\sigma\|_{\operatorname{div}_{w}}=\|\sigma\|_{\operatorname{div}}$.
The following inf-sup condition can be established in a manner similar to Theorem 3.2, and hence the proof is omitted.
Lemma 2.1.
We have the weak div stability: there exists a constant $\alpha>0$, independent of $h$, such that
(3) $\inf_{v\in\mathring{M}^{-1}_{k-1,k}}\sup_{\tau\in\Sigma_{k}^{-1}}\frac{(\operatorname{div}_{w}\tau,v)_{0,h}}{\|\tau\|_{\operatorname{div}_{w}}\|v\|_{0,h}}=\alpha>0.$
Consider the mixed formulation of the Poisson equation $\Delta u=f$ with Dirichlet boundary condition $u|_{\partial\Omega}=0$. Find $\sigma\in\Sigma_{k}^{-1}$ and $u\in\mathring{M}_{k-1,k}^{-1}$ such that
(4a) $(\sigma,\tau)$ $=(\nabla_{w}u,\tau)\quad\forall\,\tau\in\Sigma_{k}^{-1},$
(4b) $(\operatorname{div}_{w}\sigma,v)$ $=-(f,v_{0})\quad\;\forall\,v\in\mathring{M}_{k-1,k}^{-1}.$
This formulation is precisely the hybridized mixed finite element method using the ${\rm BDM}_{k}$$\mathbb{P}_{k-1}$ pair [2]. The problem (4) is well-posed by the inf-sup condition (3).
2.3. Finite element spaces
In the Stokes problem, we use the weak divergence operator to discretize the vector Laplacian. Let $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ and $\Delta\boldsymbol{u}=-\operatorname{div}\boldsymbol{\sigma}$. To distinguish the vector formulation from the scalar case, we use boldface symbols such as $\boldsymbol{\sigma}$ and $\boldsymbol{u}$.
By the tensor-product construction, we obtain the weak divergence operator
$\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{M})\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}),$
where
$\Sigma_{k}^{-1}(\mathbb{X}):=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{X})\quad\text{for }\mathbb{X}=\mathbb{M}\text{ or }\mathbb{T};\quad\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}):=\mathbb{R}^{d}\otimes\mathring{M}_{k-1,k}^{-1}.$
The tensor space is restricted to the traceless subspace $\mathbb{T}$, since for $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{div}\boldsymbol{u}=0$. We then impose suitable continuity conditions on the range space and identify the subspaces of $\Sigma_{k}^{-1}(\mathbb{T})$ that lead to weakly divergence-stable pairs.
Spaces for the velocity
We employ $H(\operatorname{div})$-conforming finite elements to discretize the velocity field $\boldsymbol{u}$. For a simplex $T$ and integers $k\geq 0$ and $\ell\geq-1$, the local space of shape functions is defined by
$\mathbb{V}_{k,\ell}^{\operatorname{div}}(T):=\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x},\quad\ell=k\text{ or }k-1.$
For $\ell\geq 0$, the degrees of freedom (DoFs) are given by (cf. [16, Section 3])
(5a) $(\boldsymbol{v}\cdot\boldsymbol{n},q)_{F},$ $\quad q\in\mathbb{P}_{k}(F),\quad F\in\mathcal{F}(T),$
(5b) $(\boldsymbol{v},\boldsymbol{q})_{T},$ $\quad\boldsymbol{q}\in{\rm grad\,}\mathbb{P}_{\ell}(T)\oplus\{\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d}):\boldsymbol{q}\cdot\boldsymbol{x}=0\}.$
When $\ell=k$, the space corresponds to the Raviart–Thomas (RT) element [47, 45], where the interior moments (5b) span $\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$. For $\ell=k-1$, the Brezzi–Douglas–Marini (BDM) element [12, 46], the interior moments (5b) forms a strict subspace.
The corresponding global finite element spaces are defined as
$\mathbb{V}_{k,\ell}^{\operatorname{div}}(\mathcal{T}_{h})$ $:=\{\boldsymbol{v}_{h}\in H(\operatorname{div},\Omega):\boldsymbol{v}_{h}|_{T}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)\text{ for all }T\in\mathcal{T}_{h}\},$
$\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}(\mathcal{T}_{h})$ $:=\mathbb{V}_{k,\ell}^{\operatorname{div}}(\mathcal{T}_{h})\cap H_{0}(\operatorname{div},\Omega).$
For simplicity, we omit $\mathcal{T}_{h}$ in the notation afterwards.
To study the weak div stability, we embed $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ into $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$.
Lemma 2.2.
The mapping $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ defined by
$\boldsymbol{u}\mapsto\big(Q_{k-1,T}\boldsymbol{u},\,(\boldsymbol{0},\,\boldsymbol{n}_{F}\cdot\boldsymbol{u})_{F}\big)_{T\in\mathcal{T}_{h},\,F\in\mathring{\mathcal{F}}_{h}}$
is injective.
Proof.
If $(\boldsymbol{n}_{F}\cdot\boldsymbol{u})|_{F}=0$ and $Q_{k-1,T}\boldsymbol{u}=0$, then all degrees of freedom in (5) vanish, which implies $\boldsymbol{u}=0$. Hence, $E$ is injective. ∎
The embedding $E$ can be extended to $H^{1}(\Omega;\mathbb{R}^{d})$, but the injectivity no longer holds.
Spaces for the pseudo-stress
For an integer $k\geq 0$, the local space of shape functions is $\mathbb{P}_{k}(T;\mathbb{T})$, with degrees of freedom (DoFs) defined by
(6a) $\int_{F}\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\,q\,\mathrm{d}S,$ $\quad q\in\mathbb{P}_{k}(F),\ i=1,2,\ldots,d-1,\ F\in\mathcal{F}(T),$
(6b) $\int_{T}\boldsymbol{\tau}:\boldsymbol{q}\,\mathrm{d}x,$ $\quad\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}).$
The unisolvence of these DoFs is established in [19, 30]. The corresponding global finite element space is
[…]
Functions in $\Sigma_{k}^{\rm tn}$ are tangential-normal continuous, i.e., $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$.
The weak divergence operator $\operatorname{div}_{w}$ defines a bilinear form on $\Sigma_{k}^{\rm tn}\times\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$:
$(\operatorname{div}_{w}\boldsymbol{\sigma},\boldsymbol{v})_{0,h}=(\operatorname{div}_{w}\boldsymbol{\sigma},E\boldsymbol{v})_{0,h}$ $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\boldsymbol{n}\cdot\boldsymbol{v})_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}\boldsymbol{v})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}])_{F}.$
Remark 2.3.
The bilinear form $(\operatorname{div}_{w}\cdot,\cdot)_{0,h}$ defines a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})^{\prime}$. With the inner product $(\cdot,\cdot)_{0,h}$, we identify $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ with its dual. Through the embedding $E$, we further identify $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cong E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})$ with its dual, and hence can write the mapping as
$\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}.$
2.4. Weak div stability
For each $T\in\mathcal{T}_{h}$, let $I_{T}^{\rm tn}:H^{1}(T;\mathbb{T})\to\mathbb{P}_{k}(T;\mathbb{T})$ denote the local interpolation operator defined by DoFs (6). The corresponding global interpolation operator $I_{k}^{\rm tn}:H^{1}(\mathcal{T}_{h};\mathbb{T})\to\Sigma_{k}^{-1}(\mathbb{T})$ is given by
$(I_{k}^{\rm tn}\boldsymbol{\tau})|_{T}:=I_{T}^{\rm tn}(\boldsymbol{\tau}|_{T}),\quad\forall\,T\in\mathcal{T}_{h},\ \boldsymbol{\tau}\in H^{1}(\mathcal{T}_{h};\mathbb{T}).$
Then $I_{k}^{\rm tn}\boldsymbol{\tau}\in\Sigma_{k}^{\rm tn}$ for $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$. For any $T\in\mathcal{T}_{h}$ and $\boldsymbol{\tau}\in H^{m}(T;\mathbb{T})$ with $1\leq m\leq k+1$, the standard interpolation estimate holds:
(7) $\|\boldsymbol{\tau}-I_{T}^{\rm tn}\boldsymbol{\tau}\|_{T}+h_{T}|\boldsymbol{\tau}-I_{T}^{\rm tn}\boldsymbol{\tau}|_{1,T}\lesssim h_{T}^{m}|\boldsymbol{\tau}|_{m,T}.$
We now prove the following important commutative property.
Lemma 2.4.
For $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$, we have
(8) $(\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}_{h})=(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h},\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
Proof.
Notice that $\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h}\in\Sigma_{k-1}^{-1}(\mathbb{T})$ and $\Pi_{F}\boldsymbol{v}_{h}]\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$. So
$(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h}$ $=-\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}I_{k}^{\rm tn}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
$=(\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}_{h}).$
This leads to the following inf-sup condition.
Theorem 2.5.
There exists a constant $\alpha>0$, independent of $h$, such that
(9) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha.$
Proof.
For any $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\subset L^{2}(\Omega;\mathbb{R}^{d})$, there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$ such that $\operatorname{div}\boldsymbol{\tau}=\boldsymbol{v}_{h}$. Let $\boldsymbol{\tau}_{h}=I_{k}^{\rm tn}\boldsymbol{\tau}$; then the commutative property (8) implies the desired result (9). ∎
If the velocity space is enriched to ${\rm RT}_{k}$, i.e., $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, then $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})$. Since DoF (6b) to define $I_{k}^{\rm tn}$ is for moments of degree k1k-1, the commutative property (8) no longer holds. Consequently, the inf-sup condition may fail for the pair $\Sigma_{k}^{\rm tn}$-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
Define the weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ as the adjoint of $\operatorname{div}_{w}$. For $\boldsymbol{v}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}\in\Sigma_{k}^{\rm tn}$ is defined by
$(\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v},\boldsymbol{\tau})=-(\boldsymbol{v},\operatorname{div}_{w}\boldsymbol{\tau})_{0,h},\quad\forall\,\boldsymbol{\tau}\in\Sigma_{k}^{\rm tn}.$
Since $\operatorname{div}_{w}\Sigma_{k}^{\rm tn}=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, the operator $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ is injective, and thus $\|\operatorname{dev}{\rm grad\,}_{w}(\cdot)\|$ defines a norm on $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$.
The inf-sup condition (9) can then be written equivalently as
(10) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|\,\|\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}_{h}\|}=1>0.$
Consider the mixed finite element method for the vector Laplacian
$\Delta\boldsymbol{u}=\boldsymbol{f},\quad\text{ with }\;\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d}).$
Find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ such that
(11) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\qquad\qquad\;\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
The formulation (11) is well posed by the inf-sup conditions (9) or (10). However, the errors $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}I^{\operatorname{div}}_{k,k-1}\boldsymbol{u}\|$ are coupled. As a result, the dominant contribution is the $H^{1}$-type error of $\boldsymbol{u}$ of order $h^{k}$, rather than the $L^{2}$ error of $\boldsymbol{\sigma}$ of order $h^{k+1}$. This suboptimality comes from the loss of exact divergence-free conformity:
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0\ \ \forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\ \ \nRightarrow\ \ (\operatorname{div}_{w}\boldsymbol{\sigma}_{h},E\boldsymbol{v})_{0,h}=0\ \ \forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
2.5. Commuting property for the weak dev grad operator
For each $T\in\mathcal{T}_{h}$, let $I_{(k,\ell),T}^{\rm div}:H^{1}(T;\mathbb{R}^{d})\to\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ denote the local interpolation operator defined by the DoFs in (5). The corresponding global interpolation operator $I_{k,\ell}^{\operatorname{div}}:H^{1}(\mathcal{T}_{h};\mathbb{R}^{d})\to L^{2}(\Omega;\mathbb{R}^{d})$ is given by
$(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})|_{T}:=I_{(k,\ell),T}^{\rm div}(\boldsymbol{v}|_{T}),\quad\forall\,T\in\mathcal{T}_{h},\ \boldsymbol{v}\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{d}).$
Then $I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}$ for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})$. For $T\in\mathcal{T}_{h}$ and $\boldsymbol{v}\in H^{m}(T;\mathbb{R}^{d})$ with $1\leq m\leq k+1$, the interpolation satisfies
(12) $\|\boldsymbol{v}-I_{(k,\ell),T}^{\rm div}\boldsymbol{v}\|_{T}+h_{T}|\boldsymbol{v}-I_{(k,\ell),T}^{\rm div}\boldsymbol{v}|_{1,T}\lesssim h_{T}^{m}|\boldsymbol{v}|_{m,T}.$
We are going to prove the following commutative property:
(13) $Q_{k}^{\rm tn}\operatorname{dev}{\rm grad\,}=\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}},$
where $Q_{k}^{\rm tn}:L^{2}(\Omega;\mathbb{T})\to\Sigma_{k}^{\rm tn}$ is the $L^{2}$-projection.
Lemma 2.6.
For any $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$,
(14) $(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})=(\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{u},\boldsymbol{\tau}_{h}),\quad\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}.$
Proof.
As $\operatorname{div}(\boldsymbol{\tau}_{h}|_{T})\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ and $\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]|_{F}\in\mathbb{P}_{k}(F)$, we have
$(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})$ $=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F},$
$=-\sum_{T\in\mathcal{T}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F}$
$=(\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{u},\boldsymbol{\tau}_{h}).$
The $\mathrm{RT}_{k}$ space is required to ensure the interior moments (5b) cover $\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ for $\operatorname{div}\boldsymbol{\tau}_{h}$, while for the $\mathrm{BDM}_{k}$ space $\mathbb{V}_{k,k-1}^{\operatorname{div}}$, the interior moments are not sufficient. Hence, the commutative property (14) does not hold for $I_{k,k-1}^{\operatorname{div}}$.
The weak operator $\operatorname{div}_{w}$ can be regarded as a discretization of $\operatorname{div}$ in the distributional sense. We can preserve the exact divergence condition in the following sense.
Lemma 2.7.
Let $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$. If
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}},$
then $\operatorname{div}\boldsymbol{\sigma}_{h}=0$.
Proof.
It suffices to prove $\operatorname{div}\boldsymbol{\sigma}_{h}=0$ in the distributional sense:
$\langle\operatorname{div}\boldsymbol{\sigma}_{h},\boldsymbol{v}\rangle=0,\quad\forall\,\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{d}).$
Let $\boldsymbol{v}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{v}$. Then by Lemma 2.6,
$\langle\operatorname{div}\boldsymbol{\sigma}_{h},\boldsymbol{v}\rangle=-(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}\boldsymbol{v})=-(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}_{I})=(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{I})_{0,h}=0.$
2.6. Restriction to divergence-free subspaces
The conditions (8) and (14) are incompatible. Weak div stability holds only on the smaller space $\mathbb{V}_{k,k-1}^{\operatorname{div}}$, while the exact divergence-freeness requires the larger space $\mathbb{V}_{k,k}^{\operatorname{div}}$. We resolve this conflict by considering the divergence-free subspaces.
We have the following commuting property (cf. [5]):
(15) $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=Q_{\ell}\big(\operatorname{div}\boldsymbol{v}\big),\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}),$
which implies $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=0$ for all $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$. The additional interior moments in ${\rm RT}_{k}$ compared with ${\rm BDM}_{k}$ serve only to enrich the range of the divergence operator. Hence, for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ and $k\geq 1$,
$I_{k,k}^{\operatorname{div}}\boldsymbol{v}=I_{k,k-1}^{\operatorname{div}}\boldsymbol{v}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
Consider the mixed finite element method for the vector Laplacian
(16) $\Delta\boldsymbol{u}=\boldsymbol{f},\quad\text{with }\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$
Find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$ such that
(17a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\qquad\qquad\;\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
(17b) $(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
Theorem 2.8.
Let $k\geq 0$. The mixed finite element method (17) is well posed. Let $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$, where $\boldsymbol{u}$ solves (16). Then
(18) $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|=\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\leq\|\boldsymbol{\sigma}-I_{k}^{\rm tn}\boldsymbol{\sigma}\|,$
and
(19) $\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|=\|Q_{k}^{\rm tn}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h})\|\leq\|\boldsymbol{\sigma}-I_{k}^{\rm tn}\boldsymbol{\sigma}\|.$
Moreover, if $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$, then
(20) $\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|+\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|\lesssim h^{k+1}\|\boldsymbol{u}\|_{k+2}.$
Proof.
The well-posedness follows from the inf-sup condition (9) and the fact
$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
Denote by $\boldsymbol{\sigma}_{I}=I_{k}^{\rm tn}\boldsymbol{\sigma}$ and $\boldsymbol{u}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$. By (8), it holds that $(\operatorname{div}_{w}\boldsymbol{\sigma}_{I},\boldsymbol{v}_{h})_{0,h}=(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v}_{h})=-(\boldsymbol{f},\boldsymbol{v}_{h})=(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$, so we conclude
$(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{v}_{h})_{0,h}=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
We verify the following orthogonality by using (14)
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})$ $=(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})+(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{h})_{0,h}$
$=(\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{I},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})=-(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{I})_{0,h}=0.$
The estimate (18) follows from this orthogonality:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|^{2}=(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}-\boldsymbol{\sigma}_{I})\leq\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{I}\|.$
By the commutative property (13), $\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{I}=Q_{k}^{\rm tn}\operatorname{dev}{\rm grad\,}\boldsymbol{u}=Q_{k}^{\rm tn}\boldsymbol{\sigma}$. Then (19) follows. ∎
The velocity space $\boldsymbol{v}_{h}|_{T}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ and $\mathbb{P}_{k}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, and the optimal order of convergence for $H^{1}$ type error is $h^{k}$. So (20) is a superconvergent error estimate for $\boldsymbol{u}$.
3. Mixed Finite Element Methods for Stokes Equation
In this section, we introduce the pressure as a Lagrange multiplier to enforce the divergence-free condition, and another multiplier to relax the tangential-normal continuity of the stress space. The resulting mixed finite element methods for the Stokes equation are pointwise divergence-free and superconvergent.
3.1. Spaces of Lagrange multipliers
The subspace $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$ is not constructed explicitly. Instead, we enforce the constraint $\operatorname{div}\boldsymbol{u}=0$ by introducing a Lagrange multiplier, which has the physical interpretation of pressure.
In the discretization, we use the piecewise polynomial space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ to approximate the pressure $p$. By the commutative property (15), it follows that
$\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}.$
We introduce the Lagrange multiplier spaces
$\Lambda_{k}=\mathbb{P}_{k}(\mathcal{F}_{h};\mathbb{R}^{d-1})\quad\textrm{and}\quad\mathring{\Lambda}_{k}=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$
to relax the tangential-normal continuity condition in the space $\Sigma_{k}^{\rm tn}$. The embedding $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ is extended as
$(\boldsymbol{u},\boldsymbol{\lambda})\mapsto\big(Q_{k-1,T}\boldsymbol{u},\,(\boldsymbol{\lambda},\,\boldsymbol{n}_{F}\cdot\boldsymbol{u})_{F}\big)_{T\in\mathcal{T}_{h},\,F\in\mathring{\mathcal{F}}_{h}},$
which remains injective.
The stress space is relaxed to $\Sigma_{k}^{-1}(\mathbb{T})=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{T})$. The weak divergence operator $\operatorname{div}_{w}$ then induces the bilinear form on $\Sigma_{k}^{-1}(\mathbb{T})\times(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$:
(21) $(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h}=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}\!\Big(([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{n}\!\cdot\!\boldsymbol{v})_{F}+([\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{\mu})_{F}\Big),$
which induces a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{T})\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})^{\prime}\cong(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$ by the $(\cdot,\cdot)_{0,h}$ inner product through the embedding $E$.
Its adjoint, still denoted by $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\Sigma_{k}^{-1}(\mathbb{T})$, is defined by
$(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu})):=-(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h},$
for $\boldsymbol{\sigma}\in\Sigma_{k}^{-1}(\mathbb{T})$, $\boldsymbol{v}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, and $\boldsymbol{\mu}\in\mathring{\Lambda}_{k}$.
The operator $\operatorname{dev}{\rm grad\,}_{w}$ can be extended to
$\operatorname{dev}{\rm grad\,}_{w}:H^{1}(\mathcal{T}_{h};\mathbb{R}^{d})\times L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})\to\Sigma_{k}^{-1}(\mathbb{T})$
as follows: for $(\boldsymbol{v},\boldsymbol{\mu})\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{d})\times L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$, let $\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu})\in\Sigma_{k}^{-1}(\mathbb{T})$ be defined elementwise by
$(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu}),\boldsymbol{\tau})_{T}=-(\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v},\,\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\,\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T},$
for all $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$.
Lemma 3.1.
For $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})$ and $\boldsymbol{\mu}\in L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$, it holds that
(22) $\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v},\,Q_{k,\mathcal{F}_{h}}\boldsymbol{\mu})=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\,\boldsymbol{\mu}),$
where $Q_{k,\mathcal{F}_{h}}$ is the $L^{2}$-orthogonal projector from $L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$ onto $\Lambda_{k}$.
Proof.
By the definition of the operator $\operatorname{dev}{\rm grad\,}_{w}$, we have for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$ that
$\quad\;(\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v},Q_{k,\mathcal{F}_{h}}\boldsymbol{\mu}),\boldsymbol{\tau})_{T}$
$=-(I_{k,k}^{\operatorname{div}}\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot(I_{k,k}^{\operatorname{div}}\boldsymbol{v}),\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}.$
Then (22) follows from the definition of the interpolation operator $I_{k,k}^{\operatorname{div}}$. ∎
3.2. Weak div stability
We first establish the following inf-sup condition.
Theorem 3.2.
There exists $\alpha>0$, independent of $h$, such that
(23) $\inf_{(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})\|_{0,h}}=\alpha.$
Proof.
Given $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, construct $\boldsymbol{\tau}_{1}\in\Sigma_{k}^{-1}(\mathbb{T})$ elementwise by
$\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}=-\tfrac{1}{2}h_{F}\,\boldsymbol{\mu}_{h}\ \ \text{on }F\in\mathcal{F}(T),\qquad(\boldsymbol{\tau}_{1},\boldsymbol{q})_{T}=0\ \ \forall\,\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}),$
for each $T\in\mathcal{T}_{h}$, where $\boldsymbol{n}$ is the unit outward normal on $\partial T$. Then $\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]=h_{F}\boldsymbol{\mu}_{h}$. By inverse and scaling estimates,
$\|\boldsymbol{\tau}_{1}\|_{\operatorname{div}_{w}}\lesssim\|h_{F}^{1/2}\boldsymbol{\mu}_{h}\|_{\mathcal{F}_{h}}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}.$
By the definition of $\operatorname{div}_{w}$,
$(\operatorname{div}_{w}\boldsymbol{\tau}_{1},(0,\boldsymbol{\mu}_{h}))_{0,h}=-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}],\boldsymbol{\mu}_{h})_{F}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}^{2},$
and hence
$\|(0,\boldsymbol{\mu}_{h})\|_{0,h}\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(0,\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}}\qquad\forall\,\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
On the other hand, by the weak divergence inf-sup (cf. (9)),
$\|(\boldsymbol{v}_{h},0)\|_{0,h}=\|\boldsymbol{v}_{h}\|\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}}\qquad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}},\ \boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
Combining the two bounds yields (23). ∎
We use $(\cdot,\cdot)_{0,h}$ as a Riesz map to identify $E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$ with its dual. The bilinear form (21) thus induces a mapping
$\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{T})\to\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
The inf-sup condition (23) implies $\operatorname{div}_{w}$ is surjective. Consequently the adjoint $\operatorname{dev}{\rm grad\,}_{w}$ is injective and
$\|\operatorname{dev}{\rm grad\,}_{w}(\cdot,\cdot)\|\text{ defines a norm on }\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
Notice that the velocity space cannot be enlarged to $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$. But later on we will use the coercivity on the divergence free subspaces.
3.3. Mixed finite element methods
With the operator $\operatorname{dev}{\rm grad\,}_{w}$, we formulate the following mixed finite element method for the Stokes equation (1): Find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}$, and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$, for integers $k\geq 0$ and $\ell\in\{k,\,k-1\}$, such that
(24a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\,\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))+(\operatorname{div}\boldsymbol{v}_{h},\,p_{h})$ $=(\boldsymbol{f},\,\boldsymbol{v}_{h}),$
(24b) $(\operatorname{div}\boldsymbol{u}_{h},\,q_{h})$ $=0,$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, and $q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$.
Theorem 3.3.
The mixed finite element method (24) is well-posed, with a unique solution $(\boldsymbol{u}_{h},\,\boldsymbol{\lambda}_{h},\,p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\times\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$. Let $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$; then $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$ satisfy the mixed formulation (17).
Proof.
We first show that the mixed method (24) admits only the zero solution when $\boldsymbol{f}=0$. From (24b), it follows that $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$. Taking $\boldsymbol{v}_{h}=\boldsymbol{u}_{h}$ and $\boldsymbol{\mu}_{h}=\boldsymbol{\lambda}_{h}$ in (24a), we obtain $\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})=0$. Combining this with (23) gives $\boldsymbol{u}_{h}=0$ and $\boldsymbol{\lambda}_{h}=0$. Furthermore, by $\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$, equation (24a) implies $p_{h}=0$. Therefore, the mixed finite element method (24) is well-posed.
Since $\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$, equation (24b) implies that $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$. Restricting $\boldsymbol{v}_{h}$ to this subspace, equation (24a) becomes
(25) $(\boldsymbol{\sigma}_{h},\,\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))=(\boldsymbol{f},\,\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}),\ \boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
Taking $\boldsymbol{v}_{h}=0$ in (25) yields $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$. Then (25) reduces to equation (17b), while equation (17a) follows directly from the definition of $\boldsymbol{\sigma}_{h}$. ∎
We circumvent the inf-sup condition, as the error analysis will be derived through the equivalence to (17) without relying on the inf-sup constant.
3.4. Error analysis
Hereafter, let $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ denote the solution of the Stokes equation (1), and set $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$. Then the first equation in (1) becomes
(26) $\operatorname{div}\boldsymbol{\sigma}-\nabla p=\boldsymbol{f}\quad\text{ in }\Omega.$
Moreover, let $\boldsymbol{\lambda}\in L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$ be defined facewise by
$\boldsymbol{\lambda}|_{F}=\Pi_{F}\boldsymbol{u},\quad\forall\,F\in\mathcal{F}_{h}.$
This notation will be used throughout the subsequent analysis.
Theorem 3.4.
Let $(\boldsymbol{u},p)$ denote the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Assume $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$. Then
(27) $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\|$
$+\|Q_{\ell}p-p_{h}\|$ $\lesssim h^{k+1}|\boldsymbol{u}|_{k+2},$
where $\|\boldsymbol{\tau}\|_{0,h}^{2}:=\|\boldsymbol{\tau}\|^{2}+\sum_{F\in\mathcal{F}_{h}}h_{F}\|\Pi_{F}\boldsymbol{\tau}\boldsymbol{n}\|_{F}^{2}$.
Proof.
By the equivalence of (24) and (17), the estimate of $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ follows from (20) in Theorem 2.8.
Then by Lemma 3.1,
$\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda})$ $=Q_{k}\operatorname{dev}{\rm grad\,}\boldsymbol{u}=Q_{k}\boldsymbol{\sigma},$
$\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ $=\boldsymbol{\sigma}_{h}.$
So
$\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\|=\|Q_{k}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\leq\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|.$
We now estimate the error of the pressure approximation. For $\boldsymbol{v}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$, by (26),
$(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})$ $=(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p)-(\boldsymbol{f},I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})+(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v},0))$
$=(\operatorname{div}\boldsymbol{\sigma},I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})+(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v},0))$
$=(I_{k}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma},{\rm grad\,}_{h}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}))\!-\!\!\!\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}(I_{k}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma})\boldsymbol{n},\Pi_{F}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}))_{\partial T}$
$\quad+(\boldsymbol{\sigma}_{h}-I_{k}^{\rm tn}\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}_{w}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v},0)).$
Then we get from (7) and (12) that
$(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}|\boldsymbol{v}|_{1}.$
Finally, by the inf-sup condition for $\operatorname{div}$ operator, we have
$\|Q_{\ell}p-p_{h}\|\lesssim\sup_{\boldsymbol{v}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})}\frac{(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})}{|\boldsymbol{v}|_{1}}\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}.$
The estimate (27) demonstrates that the mixed finite element method (24) is pressure-robust.
3.5. $L^{2}$ error estimate
The superconvergence of $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ can be obtained through the duality argument. Introduce the dual problem: Find $\tilde{\boldsymbol{u}}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ and $\tilde{p}\in L^{2}(\Omega)/\mathbb{R}$ satisfy
(28) $\begin{cases}-\Delta\tilde{\boldsymbol{u}}+\nabla\tilde{p}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}&\text{ in }\Omega,\\ \qquad\;\;\operatorname{div}\tilde{\boldsymbol{u}}=0&\text{ in }\Omega.\end{cases}$
We assume that the dual problem (28) has the $H^{2}$-regularity
(29) $\|\tilde{\boldsymbol{u}}\|_{2}\lesssim\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|.$
We refer to [29, Remark I.5.6] and [44, Section 11.5] for the $H^{2}$-regularity (29) in convex domains.
Lemma 3.5.
Let $(\boldsymbol{u},p)$ denote the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Let $\tilde{\boldsymbol{u}}$ be the solution of the dual problem (28). Assume $\operatorname{skw}{\rm grad\,}\boldsymbol{f}\in L^{2}(\Omega;\mathbb{K})$. We have
(30) $(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim\begin{cases}h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}|\tilde{\boldsymbol{u}}|_{2},&k\geq 1,\\ h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|\|\tilde{\boldsymbol{u}}\|_{1},&k=0.\end{cases}$
Proof.
It follows from (24a) with $\boldsymbol{v}_{h}=I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}}$ and $\boldsymbol{\mu}_{h}=0$, $\operatorname{div}(I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}})=0$, and (26) that
[…]
That is,
$\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}(I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}}))_{T}-\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h})\boldsymbol{n},[\![\Pi_{F}(I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}})]\!])_{F}=0.$
Then
[…]
$\quad-\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h})\boldsymbol{n},[\![\Pi_{F}(\tilde{\boldsymbol{u}}-I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}})]\!])_{F}.$
When $k\geq 1$, we get from the Cauchy-Schwarz inequality and the estimate (12) of $I_{k,k}^{\operatorname{div}}$ that
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}|\tilde{\boldsymbol{u}}|_{2}.$
Next consider case $k=0$. Applying the integration by parts to the right hand side of (31) to have
[…]
Since $\operatorname{div}\tilde{\boldsymbol{u}}=0$ and $\operatorname{div}H_{0}^{2}(\Omega;\mathbb{K})=H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ (cf. [24, Theorem 1.1]), there exists a $\tilde{\boldsymbol{\tau}}\in H_{0}^{2}(\Omega;\mathbb{K})$ such that
$\operatorname{div}\tilde{\boldsymbol{\tau}}=\tilde{\boldsymbol{u}},\quad\textrm{and}\quad\|\tilde{\boldsymbol{\tau}}\|_{2}\lesssim\|\tilde{\boldsymbol{u}}\|_{1}.$
Recall the linear finite element space of the differential (d2)(d-2)-form in skew-symmetric form [1, 5]
$\mathbb{V}_{h}^{d-2}:=\{\boldsymbol{\tau}_{h}\in H(\operatorname{div},\Omega;\mathbb{K}):\boldsymbol{\tau}_{h}|_{T}\in\mathbb{P}_{1}(T;\mathbb{K})\;\;\textrm{ for }T\in\mathcal{T}_{h}\}.$
The DoFs for space $\mathbb{V}_{h}^{d-2}$ are given in [13, (3.10)]. Let $I_{h}^{d-2}:H^{2}(\Omega;\mathbb{K})\to\mathbb{V}_{h}^{d-2}$ be the nodal interpolation operator. It holds the following commuting property:
$\operatorname{div}(I_{h}^{d-2}\boldsymbol{\tau})=I_{0,0}^{\operatorname{div}}(\operatorname{div}\boldsymbol{\tau})\quad\forall~\boldsymbol{\tau}\in H^{2}(\Omega;\mathbb{K}).$
Then, employing the integration by parts and (26),
[…]
This together with the interpolation error estimate of $I_{h}^{d-2}$ gives
$(\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\||\tilde{\boldsymbol{\tau}}|_{2}\lesssim h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|\|\tilde{\boldsymbol{u}}\|_{1}.$
This ends the proof. ∎
Theorem 3.6.
Let $(\boldsymbol{u},p)$ denote the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Assume $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$, $\operatorname{skw}{\rm grad\,}\boldsymbol{f}\in L^{2}(\Omega;\mathbb{K})$, and the $H^{2}$-regularity (29) holds. Then
(32) $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|\lesssim h^{k+2}(|\boldsymbol{u}|_{k+2}+\delta_{k0}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|),$
where $\delta_{00}=1$ and $\delta_{k0}=0$ for $k\geq 1$.
Proof.
Set $\boldsymbol{v}_{h}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$ for ease of presentation. Let $\tilde{\boldsymbol{\sigma}}={\rm grad\,}\tilde{\boldsymbol{u}}$. By the fact $\operatorname{div}\boldsymbol{v}_{h}=0$, and the definition of the operator $I_{k}^{\rm tn}$, we have
$\|\boldsymbol{v}_{h}\|^{2}$ $=-(\operatorname{div}\tilde{\boldsymbol{\sigma}},\boldsymbol{v}_{h})=\sum_{T\in\mathcal{T}_{h}}(\tilde{\boldsymbol{\sigma}},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}-\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}\tilde{\boldsymbol{\sigma}}\boldsymbol{n})_{\partial T}$
$=\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\tilde{\boldsymbol{\sigma}},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}-\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}(I_{k}^{\rm tn}\tilde{\boldsymbol{\sigma}})\boldsymbol{n})_{\partial T}$
$=\big(I_{k}^{\rm tn}\tilde{\boldsymbol{\sigma}},\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\big).$
This combined with (22) implies
[…]
Then we get from the Cauchy-Schwarz inequality, the estimate (7) of $I_{k}^{\rm tn}$, (30) and the $H^{2}$-regularity (29) that
$\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\delta_{k0}h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|.$
Thus, the estimate (32) follows from (27) and the last inequality. ∎
3.6. Postprocessing
We can construct a superconvergent approximation of $\boldsymbol{u}$ using the estimates (27) and (32). For each $T\in\mathcal{T}_{h}$, find $\boldsymbol{u}_{h}^{*}|_{T}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d})$ and $p_{h}^{*}\in\mathbb{P}_{k}(T)\cap L_{0}^{2}(T)$ such that
(33a) $(\boldsymbol{u}_{h}^{*}\cdot\boldsymbol{n},q)_{F}$ $=(\boldsymbol{u}_{h}\cdot\boldsymbol{n},q)_{F}\quad\;\;\;\forall\,q\in\mathbb{P}_{0}(F),\ F\in\mathcal{F}(T),$
(33b) $({\rm grad\,}\boldsymbol{u}_{h}^{*},{\rm grad\,}\boldsymbol{v})_{T}+(\operatorname{div}\boldsymbol{v},p_{h}^{*})_{T}$ $=(\boldsymbol{\sigma}_{h},{\rm grad\,}\boldsymbol{v})_{T}\quad\forall\,\boldsymbol{v}\in\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}),$
(33c) $(\operatorname{div}\boldsymbol{u}_{h}^{*},q)_{T}$ $=0\qquad\qquad\qquad\;\forall\,q\in{\mathbb{P}}_{k}(T)\cap L_{0}^{2}(T),$
where $\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}):=\{\boldsymbol{v}_{h}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d}):Q_{0,F}(\boldsymbol{v}_{h}\cdot\boldsymbol{n})=0\ \ \forall\,F\subset\partial T\}$.
The local problems (33) are well-posed. From (33a) we get
$Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h}^{*})=Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h})=0;$
combined with (33c), this yields $\operatorname{div}\boldsymbol{u}_{h}^{*}=0$ on each $T\in\mathcal{T}_{h}$. Thus $\boldsymbol{u}_{h}^{*}\in\mathbb{P}_{k+1}(\mathcal{T}_{h};\mathbb{R}^{d})$ is pointwise divergence-free elementwise.
Theorem 3.7.
Let $(\boldsymbol{u},p)$ be the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Assume $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$. Then
(34) $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}.$
If, in addition, $\operatorname{skw}{\rm grad\,}\boldsymbol{f}\in L^{2}(\Omega;\mathbb{K})$ and the $H^{2}$-regularity estimate (29) holds, we have
(35) $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|\lesssim h^{k+2}\big(|\boldsymbol{u}|_{k+2}+\delta_{k0}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|\big).$
Proof.
Let
$\boldsymbol{w}=(I-I_{(0,0),T}^{\operatorname{div}})(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\in\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}).$
By (15),
$\operatorname{div}\big(I_{(0,0),T}^{\operatorname{div}}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\big)=Q_{0,T}\operatorname{div}\boldsymbol{u}-Q_{0,T}\operatorname{div}\boldsymbol{u}_{h}^{*}=0.$
Hence $I_{(0,0),T}^{\operatorname{div}}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\in\mathbb{P}_{0}(T;\mathbb{R}^{d})$, and applying (15) again gives
$\operatorname{div}\boldsymbol{w}=\operatorname{div}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})=Q_{k,T}\operatorname{div}\boldsymbol{u}-\operatorname{div}\boldsymbol{u}_{h}^{*}=0.$
Using (33b) with $\boldsymbol{v}=\boldsymbol{w}$ yields
$|\boldsymbol{w}|_{1,T}^{2}$ $=({\rm grad\,}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*}),{\rm grad\,}\boldsymbol{w})_{T}$
$=({\rm grad\,}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}),{\rm grad\,}\boldsymbol{w})_{T}+(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\boldsymbol{w})_{T}.$
Applying the interpolation estimate (12) and the inverse inequality gives
(36) $\|\boldsymbol{w}\|_{T}\eqsim h_{T}|I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*}|_{1,T}\lesssim h_{T}\big(|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}|_{1,T}+\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{T}\big).$
Thus, the estimate (34) follows from the triangle inequality, the interpolation bound (12), and the stress estimate (27).
From (33a),
$I_{(0,0),T}^{\operatorname{div}}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})=I_{(0,0),T}^{\operatorname{div}}(I_{(k,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}).$
Using the triangle inequality, (12), the inverse inequality, and (36), we obtain
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|_{T}$ $\leq\|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}\|_{T}+\|I_{(0,0),T}^{\operatorname{div}}(I_{(k,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|_{T}+\|\boldsymbol{w}\|_{T}$
$\lesssim\|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}\|_{T}+\|I_{(k,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{T}$
$\quad+h_{T}\big(|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}|_{1,T}+\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{T}\big).$
Combining (12), (32), and (27) gives (35). ∎
4. Equivalent Discrete Methods
In this section, we explore several discrete methods that are equivalent to the mixed finite element scheme (24).
4.1. Hybridization of the velocity
Recall the $H^{1}$-nonconforming virtual element in [7, 15]. For integers $k\geq 0$ and $\ell=k,k-1$, the shape function space is defined as
$V_{k+1,\ell+2}^{\rm VE}(T)=\{v\in H^{1}(T):\Delta v\in\mathbb{P}_{\ell}(T),\partial_{n}v|_{F}\in\mathbb{P}_{k}(F)\textrm{ for }F\in\mathcal{F}(T)\}.$
The DoFs for $V_{k+1,\ell+2}^{\rm VE}(T)$ are given by
(37a) $(v,q)_{F},$ $\quad q\in\mathbb{P}_{k}(F),\,F\in\mathcal{F}(T),$
(37b) $(v,q)_{T},$ $\quad q\in\mathbb{P}_{\ell}(T).$
The global virtual element space $V_{k+1,\ell+2}^{\rm VE}=V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})$ by
[…]
 $\qquad\qquad\qquad\qquad\qquad\quad\quad\;\;\textrm{ single-valued across each face in $\mathring{\mathcal{F}}_{h}$}\}.$
The virtual element space $V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})$ satisfies the weak continuity condition
$([\![v]\!],q)_{F}=0\quad\forall\,v\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h}),\ q\in\mathbb{P}_{k}(F),\ F\in\mathring{\mathcal{F}}_{h}.$
When $(k,\ell)=(0,-1)$, the space $V_{1,1}^{\rm VE}$ reduces to the Crouzeix–Raviart (CR) element [25]. For $(k,\ell)=(0,0)$, the space $V_{1,2}^{\rm VE}$ corresponds to the enriched Crouzeix–Raviart element [37, 36].
Find $\boldsymbol{u}_{h}\in\prod_{T\in\mathcal{T}_{h}}\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, $\boldsymbol{\lambda}_{h}\in\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ and $p_{h}\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})/\mathbb{R}$ such that
(38a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))-(\boldsymbol{v}_{h},\nabla_{h}p_{h})$ $=(\boldsymbol{f},\boldsymbol{v}_{h}),$
(38b) $(\boldsymbol{u}_{h},\nabla_{h}q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\prod_{T\in\mathcal{T}_{h}}\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, $\boldsymbol{\mu}_{h}\in\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ and $q_{h}\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})/\mathbb{R}$.
Theorem 4.1.
The mixed method (38) is well-posed. Let $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},p_{h})\in\prod_{T\in\mathcal{T}_{h}}$ $\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ $\times\mathring{\Lambda}_{k}\times(V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})/\mathbb{R})$ be its solution, then $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},Q_{\ell}p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\times(\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R})$ is the solution of the mixed finite element method (24).
Proof.
We first prove the discrete method (38) has the zero solution when $\boldsymbol{f}=0$. By applying the integration by parts to (38b),
$\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{u}_{h},Q_{\ell,T}q_{h})_{T}-\sum_{F\in\mathcal{F}_{h}}([\![\boldsymbol{u}_{h}\cdot\boldsymbol{n}]\!],Q_{k,F}q_{h})_{F}=0\quad\forall~q_{h}\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h}).$
Thanks to DoFs (37), the above equation implies $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $\operatorname{div}\boldsymbol{u}_{h}=0$.
Let $Q_{T}^{\rm div}:L^{2}(T;\mathbb{R}^{d})\to\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ be the $L^{2}$-orthogonal projection operator. Recall the following norm equivalences on each $T\in\mathcal{T}_{h}$ established in  [39, 15, 13]:
(39) $\|Q_{T}^{\rm div}\nabla v\|_{0,T}$ $\eqsim\|\nabla v\|_{0,T}\quad\forall\ v\in V_{k+1,\ell+2}^{\rm VE}(T).$
By adding equation (38a) with $(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})=(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ and equation (38b) with $q_{h}=p_{h}$, we get $\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})=0$, which means $\boldsymbol{u}_{h}=0$ and $\boldsymbol{\lambda}_{h}=0$. Then we get from equation (38a) and the norm equivalence (39) that $p_{h}=0$. Thus, the discrete method (38) is well-posed.
For the second part, equation (38b) indicates $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $\operatorname{div}\boldsymbol{u}_{h}=0$. By restricting $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, it follows from the integration by parts that
$(\boldsymbol{v}_{h},\nabla_{h}p_{h})=(\operatorname{div}\boldsymbol{v}_{h},Q_{\ell}p_{h}).$
Hence, $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},Q_{\ell}p_{h})$ satisfies equation (24a). This ends the proof. ∎
The normal continuity for the velocity space is relaxed in (38) which is particularly useful when $\ell=-1$. As a local basis for the space $\mathring{\mathbb{V}}_{0,-1}^{\operatorname{div}}$ is not readily available, we can implement the equivalent mixed method (38) instead of the mixed method (24).
Even for $(k,\ell)=(0,-1)$, our method has the following estimates:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|+\|\operatorname{dev}{\rm grad\,}_{w}(I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{0,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\|\lesssim h|\boldsymbol{u}|_{2},$
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0}+h\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h})\|\lesssim h(|\boldsymbol{u}|_{2}+|\boldsymbol{u}|_{1}),$
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|+\|I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|\lesssim h^{2}(|\boldsymbol{u}|_{2}+\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|).$
4.2. Stabilization-free virtual element methods
For $k\geq 0$, define the space of shape functions for an $H^{1}$-nonconforming virtual element as
$\mathbb{V}_{k+1}^{\rm VE}(T):=\{\boldsymbol{v}\in H^{1}(T;\mathbb{R}^{d}):$ $\operatorname{div}\boldsymbol{v}\in\mathbb{P}_{k}(T),\textrm{ there exists some }s\in L^{2}(T)$
$\textrm{ such that }\Delta\boldsymbol{v}+\nabla s\in\big(\mathbb{P}_{k-1}(T;\mathbb{R}^{d})\cap\ker(\cdot\boldsymbol{x})\big),$
$\textrm{ and }(\partial_{n}\boldsymbol{v}+s\boldsymbol{n})|_{F}\in\mathbb{P}_{k}(F;\mathbb{R}^{d})\;\forall~F\in\mathcal{F}(T)\}.$
Following the argument in [53, Section 3.1], the local virtual element space $\mathbb{V}_{k+1}^{\rm VE}(T)$ is uniquely determined by the following DoFs
(40a) $(\boldsymbol{v},\boldsymbol{q})_{F},$ $\quad\boldsymbol{q}\in\mathbb{P}_{k}(F;\mathbb{R}^{d}),F\in\mathcal{F}(T),$
(40b) $(\boldsymbol{v},\boldsymbol{q})_{T},$ $\quad\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d}).$
Clearly we have $\mathbb{P}_{k+1}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k+1}^{\rm VE}(T)$, and $\mathbb{V}_{1}^{\rm VE}(T)=\mathbb{P}_{1}(T;\mathbb{R}^{d})$. Then define the global virtual element space by
[…]
A mixed virtual element method for the Stokes equation (1) is to find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $p_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$ with integer $k\geq 0$, such that
(41a) $(Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{u}_{h}),Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}))+(\operatorname{div}_{h}\boldsymbol{v}_{h},p_{h})$ $=(\boldsymbol{f},I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h}),$
(41b) $(\operatorname{div}_{h}\boldsymbol{u}_{h},q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $q_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$. When $k=0$, the mixed method (41) is exactly the modified Crouzeix-Raviart element method in [42, (23)].
Lemma 4.2.
It holds
(42) $\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{v}_{h}))=Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h})\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}.$
Proof.
By the definitions of the operator $\operatorname{dev}{\rm grad\,}_{w}$ and the operator $I_{k,k}^{\operatorname{div}}$, we have for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$ that
$\quad\;(\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{v}_{h})),\boldsymbol{\tau})_{T}$
$=-(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h}),\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}$
$=-(\boldsymbol{v}_{h},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v}_{h},\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}.$
Then (42) follows from the integration by parts. ∎
As a consequence of (42), $\|Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h})\|$ defines a norm on the space $\mathring{\mathbb{V}}_{h}^{\rm VE}\cap\ker(\operatorname{div}_{h})$.
Theorem 4.3.
The mixed virtual element method (41) is well-posed. Let $(\boldsymbol{u}_{h},p_{h})\in\mathring{\mathbb{V}}_{h}^{\rm VE}\times\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$ be its solution. Then $(I_{k,k}^{\operatorname{div}}\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{u}_{h}),Q_{\ell}p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h}$; $\mathbb{R}^{d-1})$ $\times\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ be the solution of the mixed finite element method (24).
Proof.
The well-posedness of the mixed finite element method (41) can be proved using an argument analogous to that in Theorem 4.1.
We then prove the equivalence between the discrete method (41) and the discrete method (24). By (42), the equation (41a) means that $(I_{k,k}^{\operatorname{div}}\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{u}_{h}),Q_{\ell}p_{h})$ satisfies equation (24a). By the commuting property (15), the equation (41a) indicates that $I_{k,k}^{\operatorname{div}}\boldsymbol{u}_{h}$ satisfies equation (24b). ∎
Remark 4.4.
We can also use the following virtual element space for the velocity:
[…]
4.3. Discrete methods based on the pseudostress-velocity-pressure formulation
By introducing $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\in\Sigma_{k}^{\rm tn}$, we can rewrite the mixed finite element method (24) as the MCS method [31]: find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$, $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ such that
$(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},q_{h};\boldsymbol{u}_{h})$ $=0$ $\forall~\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R},$
$b_{h}(\boldsymbol{\sigma}_{h},p_{h};\boldsymbol{v}_{h})$ $=-(\boldsymbol{f},\boldsymbol{v}_{h})$ $\forall~\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}},$
where the bilinear form
$b_{h}(\boldsymbol{\tau}_{h},q_{h};\boldsymbol{v}_{h})$ $:=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{T}-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n},\boldsymbol{n}\cdot\boldsymbol{v}_{h})_{\partial T}-(\operatorname{div}\boldsymbol{v}_{h},q_{h})$
$\;=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau}_{h},{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n},\Pi_{F}\boldsymbol{v}_{h})_{\partial T}-(\operatorname{div}\boldsymbol{v}_{h},q_{h}).$
The subspace
$\left\{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}:\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})\ \text{for all }F\in\mathcal{F}_{h}\right\}$
was employed in the MCS method [31] to discretize the traceless stress tensor $\boldsymbol{\sigma}$. However, the constraint $\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})$ prevents the discrete scheme in [31] from achieving superconvergence properties such as the estimates $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|Q_{\ell}p-p_{h}\|$ in (27), and $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ in (32).
Superconvergent results were later obtained in [30] by enriching the stress space, but their approach requires $k=\ell\geq 1$, whereas our formulation also covers low-order cases: $(k,\ell)=(0,-1)$, $(0,0)$, and $(1,0)$. Our approach also leads to a conceptually simpler and technically more transparent error analysis.
5. Numerical examples
In this section, we present numerical examples to validate the error estimates of the mixed finite element method (24) and the post-processing procedure (33). The test problems are adapted from [30], which primarily considered higher-order cases with $k=\ell\geq 1$. Here, we focus on low-order settings with $\ell=0$ and $k=0,1$.
The computational domain is $\Omega=(0,1)^{d}$ for $d=2,3$, uniformly partitioned into simplicial meshes. All numerical experiments were developed based on the MATLAB package iiFEM [14].
Example 5.1.
We first test the mixed finite element method (24) in two dimensions. Take the exact solution
$\boldsymbol{u}=\operatorname{curl}\psi_{2},\quad\text{and}\quad p=-x^{5}-y^{5}+\frac{1}{3},$
where $\psi_{2}=x^{2}(x-1)^{2}y^{2}(y-1)^{2}$.
Table 1. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ order $\|p-p_{h}\|$ order
$2^{-3}$ $(0,0)$ 2.988e-03 3.103e-02 7.810e-02
$2^{-4}$ $(0,0)$ 1.284e-03 1.22 1.677e-02 0.89 3.914e-02 1.00
$2^{-5}$ $(0,0)$ 5.988e-04 1.10 8.700e-03 0.95 1.963e-02 1.00
$2^{-6}$ $(0,0)$ 2.907e-04 1.04 4.440e-03 0.97 9.840e-03 1.00
$2^{-7}$ $(0,0)$ 1.436e-04 1.02 2.247e-03 0.98 4.931e-03 1.00
$2^{-3}$ $(1,0)$ 3.296e-04 2.447e-03 7.453e-02
$2^{-4}$ $(1,0)$ 8.382e-05 1.98 6.305e-04 1.96 3.760e-02 0.99
$2^{-5}$ $(1,0)$ 2.104e-05 1.99 1.597e-04 1.98 1.880e-02 1.00
$2^{-6}$ $(1,0)$ 5.266e-06 2.00 4.016e-05 1.99 9.428e-03 1.00
$2^{-7}$ $(1,0)$ 1.317e-06 2.00 1.007e-05 2.00 4.715e-03 1.00
Table 2. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-3}$ $(0,0)$ 1.233e-03 2.890e-02
$2^{-4}$ $(0,0)$ 3.277e-04 1.91 1.481e-02 0.96
$2^{-5}$ $(0,0)$ 8.353e-05 1.97 7.453e-03 0.99
$2^{-6}$ $(0,0)$ 2.099e-05 1.99 3.733e-03 1.00
$2^{-7}$ $(0,0)$ 5.256e-06 2.00 1.867e-03 1.00
$2^{-3}$ $(1,0)$ 3.296e-05 2.286e-03
$2^{-4}$ $(1,0)$ 4.167e-06 2.98 5.183e-04 1.98
$2^{-5}$ $(1,0)$ 5.264e-07 2.99 1.463e-04 1.99
$2^{-6}$ $(1,0)$ 6.625e-08 2.99 3.666e-05 2.00
$2^{-7}$ $(1,0)$ 8.315e-09 2.99 9.178e-06 2.00
Example 5.2.
We then test the mixed finite element method (24) in three dimensions. Take the exact solution
$\boldsymbol{u}=\operatorname{curl}(\psi_{3},\psi_{3},\psi_{3})^{\intercal},\quad\text{and}\quad p=-x^{5}-y^{5}-z^{5}+\frac{1}{2},$
where $\psi_{3}=x^{2}(x-1)^{2}y^{2}(y-1)^{2}z^{2}(z-1)^{2}$.
The numerical errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ are reported in Table 1 for Example 5.1 and Table 3 for Example 5.2, while the post-processed errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ are given in Table 2 for Example 5.1 and Table 4 for Example 5.2. The results in these tables demonstrate that
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|\eqsim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h^{k+1}),\quad\|p-p_{h}\|=\mathcal{O}(h),$
and for the post-processed solution $\boldsymbol{u}_{h}^{*}$,
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|=\mathcal{O}(h^{k+2}),\quad\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|=\mathcal{O}(h^{k+1}),$
which are consistent with the theoretical estimates (27) and (34)–(35).
Table 3. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$, and $\|p-p_{h}\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order
$2^{-1}$ $(0,0)$ 6.762e-04 9.787e-03 2.942e-01
$2^{-2}$ $(0,0)$ 3.110e-04 1.12 4.960e-03 1.06 1.649e-01 0.84
$2^{-3}$ $(0,0)$ 1.399e-04 1.15 2.431e-03 0.95 8.501e-02 0.96
$2^{-4}$ $(0,0)$ 6.719e-05 1.06 1.260e-03 0.95 4.284e-02 0.99
$2^{-5}$ $(0,0)$ 3.320e-05 1.02 6.397e-04 0.98 2.146e-02 1.00
$2^{-1}$ $(1,0)$ 2.076e-04 2.431e-03 2.942e-01
$2^{-2}$ $(1,0)$ 6.700e-05 1.63 5.715e-04 2.09 1.649e-01 0.84
$2^{-3}$ $(1,0)$ 1.790e-05 1.90 1.541e-04 1.89 8.501e-02 0.96
$2^{-4}$ $(1,0)$ 4.551e-06 1.98 3.975e-05 1.96 4.284e-02 0.99
Table 4. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-1}$ $(0,0)$ 4.167e-04 4.576e-03
$2^{-2}$ $(0,0)$ 1.565e-04 1.41 2.880e-03 0.67
$2^{-3}$ $(0,0)$ 4.400e-05 1.83 1.539e-03 0.90
$2^{-4}$ $(0,0)$ 1.143e-05 1.95 7.836e-04 0.97
$2^{-5}$ $(0,0)$ 2.889e-06 1.98 3.937e-04 0.99
$2^{-1}$ $(1,0)$ 9.624e-05 1.894e-03
$2^{-2}$ $(1,0)$ 1.493e-05 2.69 5.478e-04 1.79
$2^{-3}$ $(1,0)$ 2.035e-06 2.88 1.486e-04 1.88
$2^{-4}$ $(1,0)$ 2.627e-07 2.95 3.808e-05 1.96